1) Mike is thinking about going to Australia for vacation, how many AUD will he receive if he converts $1,000 USD, without a currency exchange fee?

2) Mike is also thinking of taking a trip to Brazil, how many BRL will he receive if he converts $1,000 USD, without a currency exchange fee?

3) Instead of taking two trips Mike is now considering taking one long trip where he will go from the USA to Brazil and then to Australia. If he starts with $2,000 USD, how many BRL will he receive, without a currency exchange fee? While in Brazil Mike spends 1000 BRL, how many AUD will he receive when he converts his money into AUD, without a currency exchange fee?  If Mike spends 1000 AUD while in Australia, how many USD will he receive when he converts his money back in the USA, without a currency exchange fee?


exchange rates are suppose to be collected from

xe.com
Answer #1

Currently as per xe.com, below are the Exchange Rates.

1 USD = 1.47301 AUD

1 USD = 4.19273 BRL

1 AUD = 2.84636 BRL

1. He will receive 1000* 1.47301 = 1473.01 AUD.

2. He will receive 1000* 4.19273 = 4192.73 BRL.

3. Total BRL he will receive is 2000* 4.19273 = 8385.46 BRL.

In this, he spends 1000 BRL, then remaining he has 7385.46 BRL. If he converts into AUD, he will receive 2594.70 AUD (7385.46/2.84636).

In this, he spends 1000 AUD, then remaining he has 1594.70 AUD. If he converts into USD, he will receive 1082.61 USD (1594.70/1.47301).

Note: All values have been rounded off to 2 decimals.